In order to prepare a standard solution of 0.100M sodium carbonate, (MM:106g mol-1), 10.6g of Na2CO3 should be dissolved in:
A. 1000cm3 of water
B. enough water to make 1000cm3 of solution
C. 1000g of water
D. enough water to make 1000g of solution
1
Expert's answer
2020-07-27T07:37:01-0400

molarity=mol/l


m(Na2CO3)=10.6g

MM(Na2CO3)=106g*mol-1


n=m/M=10.6/106=0.1mol

l=n/molarity=0.1/0.1=1l=1000cm3


Answer: A. 1000 сm3 of water